1.     Locate all the instantaneous centres of the mechanism as shown in Fig.. The lengths of various links are: AB = 150 mm ; BC = 300 mm ; CD = 225 mm ; and CE = 500 mm. When the crank AB rotates in the anticlockwise direction at a uniform speed of 240 r.p.m. ; find 1. Velocity of the slider E, and 2.Angular velocity of the links BC and CE.


 



1
Expert's answer
2021-05-13T07:19:24-0400






"N=240r.p.m"

"\omega_2=\omega_{AB}=\frac{2 \pi N }{60}=25.13rad/sec"

"V_B=\omega _{AB} \times AB =25.13 \times 0.15 =3.769 m/s"

V.D Scale "25.13 m\/s =1.6m\/s"

V.D Scale "3.769 rad\/s =2.4 rad\/s"